OpenStudy (anonymous):

How would you solve k squared = 16? This is solving quadratic equations by taking square roots.

OpenStudy (anonymous):

\[k^2=16\] move the 16 over \[k^2-16=0\] this is a difference of squares

OpenStudy (anonymous):

Am i now adding 16?

OpenStudy (anonymous):

no we just moved it over difference of squares x^2-a^2=(x-a)(x+a)

OpenStudy (anonymous):

woah..

OpenStudy (anonymous):

so what exactly is the next step? o.o

OpenStudy (anonymous):

well k^2-16= (k-4)(k+4)=0 this means that k-4=0 & k+4=0 so we solve for both k's

OpenStudy (anonymous):

oh yeah!

OpenStudy (anonymous):

or you can just find the square root of both sides of the initial problem \[\sqrt{k^2}=\sqrt{16}\] just remember \[\sqrt{1}=\pm 1\]

OpenStudy (anonymous):

oh yeahhh, ok cool

OpenStudy (anonymous):

2 method of solving, both give the same answer

OpenStudy (anonymous):

so i leave the answer as k-4 =0 and k+4+0?

OpenStudy (anonymous):

*=